OpenStudy (anonymous):

Lola’s car’s gas tank is 7/8 full. Would she have more gas if she had 4/5 of a tank? Why?

OpenStudy (akshay_budhkar):

if you notice \[7/8 > 4/5\] So the tank is more filled when it is 7/8th full

OpenStudy (anonymous):

In these kind of problems we want to write them with the same denominator. What we do is that we find the greatest common denominator (GCD). In this case 5*8 is the (GCD) 7*5 / 8*5 = 35 / 40 4*8 / 5*8 = 32 / 40 35 > 32 and therefore 7/8 > 4/5

OpenStudy (anonymous):

I'm sorry 1 is the GCD and therefore you multiply them. The answer is still right though.